OpenStudy (anonymous):

Few questions 1) A car gets 30 miles per gallon of gas and is driven 1000 miles per month. Gas costs $4 per gallon. How much will the car's owner save on gas expenses in one year by buying a new car that goes 50 miles for each gallon of gas? 2) Find the value of x for which x degrees Fahrenheit is exactly the same as x degrees Celsius (remember the formula for converting Celsius to Fahrenheit is to multiply the Celsius temperature by 9/5 and add 32). 3) Simplify: [\frac{1}{1+\frac{1}{1+\frac{1}{1+\frac{1}{1+1}}}}] 4) Compute A divided by B where A = 66666^4 and B = 22222^4.

OpenStudy (tkhunny):

I'm sure answers would be welcome, until it came to your examination and you have no idea how to proceed because you always had someone else do your work for you. SHOW YOUR WORK!

OpenStudy (anonymous):

Sorry, I already tried and did quite a few other problems. I've been working all day on others like these and I am going a bit brain-dead, and I cannot figure out how to start these few. fine, I'll try the last few (again) Again, sorry but thanks

OpenStudy (tkhunny):

\(\dfrac{1\;gal}{30\;mi}\cdot\dfrac{1000\;mi}{1\;month}\cdot\dfrac{4\;dollar}{1\;gal} = 133\;1/3\;dollar/month\) \(\dfrac{1\;gal}{50\;mi}\cdot\dfrac{1000\;mi}{1\;month}\cdot\dfrac{4\;dollar}{1\;gal} = 80\;dollar/month\) A little dimensional analysis!